### v2

# Theme-Explanation Structure for Table Summarization using Large Language Models: A Case Study on Korean Tabular Data

Abstract
This paper proposes the Theme-Explanation Structure-based Table Summarization (Tabular-TX) pipeline designed to process tabular data efficiently. Tabular-TX preprocesses tabular data by focusing on highlighted cells. It then generates summary sentences following a structured format, where the Theme Part appears as an adverbial phrase, and the Explanation Part follows as a predictive clause. This approach enables tailored analysis by considering the structural characteristics of tables and their comparability. Unlike conventional fine-tuning approaches that require extensive labeled data and computational resources, our method leverages In-Context Learning to dynamically adapt to different table structures without additional training, ensuring efficient and scalable table interpretation. Experimental results demonstrate that Tabular-TX significantly outperforms conventional fine-tuning-based methods, particularly in low-resource scenarios, by leveraging table structures and metadata more effectively through structured prompts. The results confirm that Tabular-TX enables more effective processing of complex tabular data. Furthermore, it serves as a viable alternative for table-based question answering and summarization tasks in resource-constrained environments.
1 Introduction
footnotetext: * These authors contributed equally.footnotetext: $\dagger$ Corresponding author.
Tables are an essential visualization tool for concisely organizing and comparing complex data. However, tabular data is inherently compact, requiring higher analytical capabilities compared to text data. Current large language models (LLMs) face challenges in effectively handling the structural complexity of tabular data Liu et al. (2024). Therefore, enabling LLMs to process table data efficiently and perform Table Question-Answering (TableQA) remains a significant challenge.
Previous studies on TableQA primarily aimed to enhance LLM performance through fine-tuning methods Zhang et al. (2024b). However, fine-tuning requires substantial computational resources Strubell et al. (2019), and Korean tabular datasets are limited in size, making it challenging to fine-tune LLMs effectively. For example, studies leveraging table-related datasets such as FeTaQA Nan et al. (2022) have shown performance improvements through fine-tuning but suffer from high computational costs. Meanwhile, synthetic dataset generation techniques Hou et al. (2024) have been proposed to address data scarcity. However, repetitive training on generated data has been shown to degrade model performance Alemohammad et al. (2024).
This study aims to enhance LLMs’ in-context learning (ICL) ability to generate interpretations of table segments using the table interpretation corpus provided by the National Institute of Korean Language NIKL (2024). As shown in Figure 2, the NIKL task is to summarize the highlighted cells, which are labeled in the metadata as highlighted_cells, into a single coherent sentence. To achieve this, we propose the Theme-Explanation Structure-based Table Summarization (Tabular-TX) pipeline.
Tabular-TX simplifies the complexity of tabular data and conducts tailored analysis considering its structural characteristics. Specifically, it segments table summaries into two distinct components: the Theme Part, expressed as an adverbial phrase, and the Explanation Part, presented as a predicative clause. This structured approach ensures a more comprehensive analysis by considering the metadata and the cell data. Furthermore, we introduce a journalist persona into the LLMs to enhance the objectivity and clarity of the generated sentence.
Figure 1: An overall pipeline of Theme-Explanation Structure-based Table Summarization (Tabular-TX).
Our main contributions are as follows:
  • We propose the Tabular-TX, a method for generating table summaries based on the theme-explanation structure, significantly improving the interpretative performance of table segments.
  • We suggest preprocessing techniques to address complex structural issues in tabular data to improve LLM’s ability to interpret table data.
  • Our experimental results of Tabular-TX show high performance without fine-tuning by utilizing LLMs’ in-context learning (ICL) ability.
2 Related Work
2.1 Table Question Answering (TableQA)
To enable complex reasoning over tabular data, Wang et al. (2024) proposed the Chain-of-Table framework as an extension of the text-based Chain-of-Thought method Wei et al. (2022). This approach simplifies inference by reordering, extracting, and filtering table data, ultimately integrating relevant information into a structured table format. While this method excels in structured table processing and mathematical reasoning, it has limitations in generating interpretations for sections that require metadata or background knowledge.
Figure 2: An example from the corpus for evaluating interpretation generation of table segments (originally in Korean, translated into English) NIKL (2024).
TableLlama Zhang et al. (2024b) aims to generalize table-based models beyond task-specific constraints by fine-tuning 14 datasets across 11 tasks, including Highlighted Cells QA. The model achieved performance comparable to or surpassing task-specific models and even outperformed GPT-4 OpenAI (2023) on unseen tasks. However, despite its effectiveness, the model suffers from high computational costs. The training process demands 48 NVIDIA A100 80GB GPUs, making fine-tuning infeasible in resource-constrained environments.
Among table-based interpretation benchmarks, FeTaQA Nan et al. (2022) serves as a key reference dataset for evaluating table summarization models. While Chain-of-Table and TableLlama utilize ICL Brown et al. (2020) and fine-tuning, respectively, they struggle to incorporate metadata into their interpretations effectively.
2.2 Data Synthesis and In-Context Learning for Addressing Data Scarcity
Various approaches have been proposed to address the scarcity of table data. Hou et al. (2024) suggested generating synthetic table datasets through AI-driven augmentation techniques. Meanwhile, Wang et al. (2024) introduced Zero-shot Prompting and Few-shot Prompting as efficient alternatives for adapting LLMs to table tasks with minimal data. However, studies have shown that repeated training on synthetic datasets can degrade the quality and diversity of the data, leading to performance deterioration Alemohammad et al. (2024).
Few-shot prompting is a method where examples, explanations, and instructions are provided as input to guide the model’s responses without directly updating its parameters Brown et al. (2020). This technique eliminates the need for retraining the model, making it particularly useful in environments with limited datasets. In this study, we employ this approach to generate interpretations for segments of tables while minimizing computational costs.
3 Tabular-TX Pipeline
3.1 Theme-Explanation Structure
We propose Tabular-TX, a pipeline for Theme-Explanation structure-based table summarization, designed to generate interpretations for table segments. Unlike conventional approaches that treat table summaries as isolated text generation tasks, our method ensures structural consistency by explicitly organizing content into a Theme Part and an Explanation Part.
3.1.1 Theme Part
Figure 3: A sentence including the table title conveys the context more accurately. (Gray text indicates the English translation.)
The Theme Part serves as a crucial contextual anchor, ensuring that numerical or categorical values in the table are interpreted correctly. It is structured as an adverbial phrase, combining the noun phrase of the table title (table_title) with a citation or basis expression. This structure is essential because the table title provides the sole comprehensive context in table summaries. Unlike general text summaries, table cells alone are insufficient to provide meaningful context, making the resulting sentence ambiguous without additional background information.
For example, in the sentence: “According to the refugee status by nationality, the total number of refugee applications is 2,437, and among them, only 147 have been approved, indicating a very low approval rate.” Here, the Theme Part is: “According to the refugee status by nationality,” This phrase, introduced with the citation expression “According to”, provides essential context for the numerical values that follow. Without this structured introduction, the reader may struggle to understand the significance of the numbers. Figure 3 illustrates how omitting the Theme Part results in an unclear or misleading summary.
3.1.2 Explanation Part
Following the Theme Part, the Explanation Part delivers a structured analysis of the highlighted cells, forming the core content of the summary. Depending on the data type, this section uses a specific analytical technique, such as enumeration, magnitude comparison, or trend analysis. The choice of method is determined based on the comparability of the highlighted cells, ensuring that the summary provides meaningful insights rather than just raw cell values.
For instance, in the previous example, the Explanation Part is: “the net fiscal cost increased by 9.435 trillion KRW from the previous year, reaching a total of 61.301 trillion KRW.” Here, the Explanation Part is derived by comparing the numerical changes between the two cells. Here, a trend analysis is applied to highlight the increase in fiscal cost.
3.2 Sentence Generation Process
Generating sentences with a theme-explanation structure in Tabular-TX involves multiple processing steps to transform tabular data into structured natural language summaries.
3.2.1 Data Transformation
Figure 4: An example of inferring relationships between data sharing the same row or column. Through inference in ①, it is deduced that ‘10021’ represents the ‘Exam number.’ In ②, the meaning of ‘pass’ is inferred. In ③, it is deduced that the exam with the ‘Exam number’ ‘10021’ has ‘passed.’ (Gray text indicates the English translation.)
Figure 5: An example of how merged cells hinder table recognition. (Gray text indicates the English translation.)
The first step is preprocessing the table to simplify its structure for better LLM comprehension. Since LLMs primarily operate on sequential text representations, directly processing raw tabular formats can lead to misinterpretation of hierarchical relationships within the data. To address this, we convert table data into a key-value pair dictionary format, which is commonly used in natural language processing tasks. This transformation significantly enhances LLMs’ ability to recognize table semantics, improving summarization accuracy Stengel-Eskin et al. (2021).
Then, we process merged cells to clarify the table structure. Merged cells span multiple rows or columns and are defined by ‘rowspan’ and ‘colspan.’ As shown in Figure 4, LLMs infer relationships between data through row or column alignment. However, incorrect handling of merged cell ranges can lead to misinterpretation. For example, in Figure 5, the cell labeled “2020” should cover columns 3 and 4, but it appears only in column 3. To resolve this, merged cells are replicated across their ranges, allowing LLMs to recognize cell dependencies and hierarchical structures correctly.
Finally, the transformed dictionary list retains only the highlighted and related cells, where “related cells” refer to all header cells sharing the same row/column as the highlighted cells. This process reduces data complexity and enhances LLMs’ recognition of table structures.
3.2.2 Chain-of-Thought (CoT)
After data transformation, the process moves to generate interpretations for table segments using Chain-of-Thought (CoT) reasoning. This technique is particularly effective for table summarization as it systematically decomposes multi-step reasoning tasks, ensuring that relationships between numerical values and categorical data are captured accurately.
One major issue in table summarization is the Compositional Deficiency problem Zhao et al. (2024), where individual data points are analyzed separately without being adequately integrated into a holistic interpretation. Even if an LLM correctly identifies each table cell individually, the final summary may still lack contextual completeness if the relationships among the values are not recognized.
To mitigate this issue, CoT guides LLMs through a structured reasoning process that progressively refines the interpretation of highlighted table segments. First, the model classifies the type of information contained in the highlighted cells, distinguishing between monetary values, percentages, categorical data, or textual explanations. This step prevents errors, such as misinterpreting percentages as simple numbers, which could yield incorrect results during comparative analysis.
Based on the classification, it selects the most appropriate analytical method, such as enumeration for independent values, magnitude comparison for ranking numerical data, or trend analysis for detecting changes over time. For example, when processing monetary data, the model ensures proper unit conversion to maintain consistency across the summary. Similarly, when dealing with percentage values, a decimal-to-percentage transformation is applied to improve clarity. In cases where the highlighted cells contain free-text explanations, the model leverages background knowledge to provide a more context-aware interpretation.
Afterward, the most suitable analytical method for explaining the relationships between the data is selected and applied to the Explanation Part. For example, enumeration is used if the data items are individually critical, magnitude comparison is used to rank by value, and trend analysis is chosen to track changes over time. These decisions impact the summary’s informativeness.
By integrating these steps logically, CoT ensures that the final summary accurately captures relationships between data points while maintaining structural coherence and readability.
3.2.3 Journalist Persona for Structured Generation
Figure 6: An example of how a persona-applied prompt generates more accurate summaries. (Gray text indicates the English translation.)
Through strategic prompt engineering, we assign a journalist persona to the LLM to generate Theme-Explanation structured summaries. This persona is particularly effective because table summaries share key characteristics with straight news articles, which prioritize conciseness, objectivity, and fact-based clarity. Rather than generating overly detailed or speculative content, the model produces well-structured and neutral summaries that adhere to journalistic reporting conventions when guided by this persona.
Figure 6 demonstrates the impact of this persona on table summarization. When a generic prompt is used, the model generates an unstructured and somewhat ambiguous summary that captures the core information but lacks contextual clarity and fails to integrate key elements cohesively. Even with specific instructions, the generic prompt still lacks fluency and focuses solely on the facts. In contrast, with the journalist persona applied, the model produces a structured, precise, and contextually enriched summary. This transformation occurs because the journalist persona naturally guides the model to state the source of information explicitly, clearly define numerical constraints, and incorporate additional contextual details that enhance readability. Instead of merely listing numerical values, the model provides a structured sentence that mirrors how financial or administrative data would be reported in news articles.
Model ROUGE-1 ROUGE-L BLEU Average
kobart-base-v2 - Fine-tuned 0.37 0.28 0.35 0.33
EXAONE 3.0 7.8B - ICL 0.21 0.14 0.01 0.12
EXAONE 3.0 7.8B - LoRA 0.27 0.21 0.05 0.17
EXAONE 3.0 7.8B - Tabular-TX 0.51 0.39 0.44 0.45
llama-3-Korean-Bllossom-8B - ICL 0.33 0.25 0.27 0.28
llama-3-Korean-Bllossom-8B - Tabular-TX 0.48 0.37 0.42 0.43
Table 1: Evaluation metrics scores for each model.
4 Experimental Setup
4.1 Dataset and Evaluation Metrics
For training and evaluation, we utilized the table interpretation corpus provided by the National Institute of Korean Language NIKL (2024), which focuses on summarizing highlighted table segments into coherent sentences. An example of the dataset is shown in Figure 2. The dataset consists of 7,170 training tables, 876 validation tables, and 876 test tables. Each data point contains metadata such as the document title, table title, publication date, publishing organization, table source URL, highlighted cell information, table data, and a reference summary sentence describing the highlighted portions’ key contents.
We employ ROUGE-1, ROUGE-L, and BLEU to evaluate the performance of table segment interpretation. These metrics assess how effectively the summaries convey the key content of the table while achieving high semantic quality.
4.2 Models
4.2.1 Tabular-TX Pipeline Application
To evaluate the effectiveness of the Tabular-TX pipeline, We utilize EXAONE 3.0 7.8B An et al. (2024) and llama-3-Korean-Bllossom-8B [^1] https://huggingface.co/MLP-KTLim/llama-3-Korean-Bllossom-8B models as base models. EXAONE 3.0 7.8B, a successor to EXAONE-LM-v1.0, has demonstrated state-of-the-art performance in Korean TableQA, ranking first on the KorWikiTableQuestions Jun et al. (2022). Similarly, llama-3-Korean-Bllossom-8B is the top-performing sub-10B model in a Korean multi-domain reasoning benchmark. We compare the performance of these models with and without Tabular-TX, assessing whether structured generation (Theme-Explanation format) enhances performance in table summarization. Additionally, we analyze whether Tabular-TX reduces reliance on extensive fine-tuning while maintaining high-quality summaries.
4.2.2 Additional Adaptation Approaches
In-Context Learning
We begin by applying In-Context Learning (ICL) to each model, providing a few table-summarization examples without explicit fine-tuning. This approach tests how effectively the model can generate coherent sentences for highlighted table cells based solely on a small set of demonstrations.
Low-Rank Adaptation
Next, we assess the computational efficiency and performance of the Tabular-TX pipeline by introducing LoRA (Low-Rank Adaptation). We apply LoRA to the EXAONE 3.0 7.8B model to see if we can maintain high-quality table summaries with fewer resources.
Full Model Fine-Tuning
Finally, we evaluate the gogamza/kobart-base-v2 model under a full model fine-tuning setup to determine whether a smaller-scale language model can achieve comparable performance when all its parameters are trained on the table interpretation corpus.
5 Experimental Results
Table 1 presents the performance of various models evaluated using ROUGE-1, ROUGE-L, BLEU, and their average scores. The kobart-base-v2 recorded an average score of 0.33 after fine-tuning. In contrast, EXAONE 3.0 7.8B achieved 0.12 with the ICL method, 0.17 after fine-tuning, and 0.45 when combined with the Tabular-TX method. Similarly, llama-3-Korean-Bllossom-8B, which was also tested with Tabular-TX, showed a notable improvement, reaching an average score of 0.43. These results demonstrate that Tabular-TX consistently outperforms alternative methods, achieving the highest overall performance across different model configurations.
The performance gap between EXAONE 3.0 7.8B and kobart-base-v2, despite both being fine-tuned on the same dataset, can be explained through the multiplicative joint scaling law Zhang et al. (2024a). This principle suggests that when the dataset size is insufficient relative to the model size, the performance gains from fine-tuning remain limited. Since kobart-base-v2 has 124M parameters, while EXAONE 3.0 is approximately 63 times larger, the dataset required to achieve a comparable performance boost must be proportionally scaled up by a factor of 63. The inability to meet this scaling requirement explains why kobart-base-v2’s fine-tuned performance plateaued, while EXAONE 3.0 7.8B demonstrated more significant gains with the same dataset.
This study confirms that the proposed Tabular-TX method enhances table data analysis performance without fine-tuning. Notably, Tabular-TX outperforms traditional fine-tuned models despite relying on significantly smaller datasets, demonstrating its efficacy in resource-constrained learning environments. Moreover, Tabular-TX achieved approximately four times higher average performance compared to standard ICL methods, further reinforcing its role as a scalable and efficient alternative for structured table summarization tasks.
6 Conclusion
This study introduced the Theme-Explanation Table Summarization (Tabular-TX) pipeline, a novel approach to improve table summarization tasks with low-resource requirements.
Experimental results signifies that Tabular-TX enhances table summarization performance. This study contributed to the summarization of complex table data by introducing a novel sentence generation method based on the theme-explanation structure. Furthermore, Tabular-TX achieved excellent performance without fine-tuning, by incorporating ICL. This indicates its potential as a significant contribution to table data analysis, even in resource-constrained environments, without requiring direct model training.
Limitations
We acknowledge a few limitations in this study. First, Tabular-TX was only evaluated on EXAONE 3.0 7.8B and llama-3-Korean-Bllossom-8B, leaving the question of its effectiveness across a broader range of LLMs open. Second, this study primarily focused on Korean administrative table data, and further research should investigate whether the Theme-Explanation structure is equally effective for diverse tabular data formats in other languages or specialized domains. Finally, Tabular-TX currently relies on predefined structural components (Theme and Explanation parts) to enforce interpretability. Future work should explore more dynamic approaches that allow for adaptive sentence structuring based on different types of tables, potentially improving performance across varied tabular structures.
Appendix A Tabular-TX Construction Details
Tabular-TX generates table summaries in two steps. First, the Data Recognition / Classification step identifies key data from highlighted cells (Figure 7). Second, the Sentence Generation step forms a summary in the Theme-Explanation format (Figure 8).
Figure 7: Summarizing key data points from the table in a single sentence for a news article. (originally in Korean, translated into English)
Figure 8: Writing a one-sentence summary of a table by embodying news reporter persona. (originally in Korean, translated into English)